The length of the base of a rriangle is twice its height. If the triangle is 100 square kilometers, find the height
Answer by umnikmo(75)

S=1/2*b*h
b=2*h
S=1/2*2*h*h=h^2=100 => h=sqrt(100) = 10
b=2*10=20